Find an equation of the line containing the given pair of points (4,5) and (12,9)
Answer by algebrahouse.com(1659)

y2 - y1
-------- = slope
x2 - x1

(4,5) , (12,9)

9 - 5
------ = slope {substituted coordinates into slope formula}
12 - 4

4
-- = slope {subtracted}
8

1
-- = slope {reduced}
2


(4,5) , m = 1/2 {point and slope}
y - y1 = m(x - x1) {point-slope formula}

y - 5 = (1/2)(x - 4) {substituted point and slope into point-slope formula}
y - 5 = (1/2)x - 2 {used distributive property}
y = (1/2)x + 3 {added 5 to each side}

For more help from me, visit: www.algebrahouse.com